1.We have a Poisson distribution with mean = 4. Find P(X = 3), P(X < 3), P(X <= 3), P(X > 3), the variance, and standard deviation.


2.Three members of a club will be selected to serve as officers. The first person selected will be president, the second person will be vice-president and the third will be secretary/treasurer. How many ways can these officers be selected if there are 30 club members?
A) 90
B) 27000
C) 52360
D) 4060
E) 24,360

Answer by stanbon(75887)

Three members of a club will be selected to serve as officers. The first person selected will be president, the second person will be vice-president and the third will be secretary/treasurer. How many ways can these officers be selected if there are 30 club members?
---
# of ways to select a president: 30
# of ways to select a vp: 29
# of ways to select sec/tres: 28
-----
Totoal # of choices: 30*29*28 = 24360
---
This is the same as 30P3 = 30!/(30-3)! = 30*29*28
You are choosing 3 from 30 and order makes a difference.
=====================================
Cheers,
Stan H.
=======================================
A) 90
B) 27000
C) 52360
D) 4060
E) 24,360